OpenStudy (anonymous):

MEDAL GIVEN!! the sum of a square number and five times the number itself is 24. By solving a quadratic equation, find the 2 possible values of the number.

OpenStudy (anonymous):

@Abdulhameed

OpenStudy (anonymous):

first u have to make this question in the form of equation

OpenStudy (anonymous):

5x^2-24=0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

let the number is x then square the number is x^2 and five times the number is 5x

OpenStudy (anonymous):

then x^2+5x=24 x^2+5x-24=0 solve the equation

OpenStudy (anonymous):

@ryanvarghese12790

OpenStudy (anonymous):

ok x=-8 or x = 3

OpenStudy (anonymous):

yeah ur solution is correct :)